OpenStudy (anonymous):

Find 4x − 2y − z, if x = 2, y = 3, and z = −6 please help !

OpenStudy (anonymous):

4(2) - 2(3) +6 = 8

OpenStudy (anonymous):

4(2)-2(3)-(-6) =8-6+6 =4

OpenStudy (anonymous):

oops haha =8